Information-centric collaborative sites have emerged as viable sources for many information needs. In particular, the Community Question Answering (CQA) model has evolved as an effective alternative to web search: users post questions and answers, and rate and evaluate each others content. As a result, users contribute both content and rich interaction metadata in their roles as askers, answerers, and evaluators. To give a sense of the scale, just one popular site, Yahoo! Answers, has attracted an active participation of millions of users, and has amassed a searchable archive of hundreds of millions of answers to tens of millions of questions. The talk will build on our earlier work on modeling user interactions in web search, and will describe new problems related to mining this user-generated content, including estimating content quality, contributor authority, user intent, and asker satisfaction with the results.